OpenStudy (kimberlyevens):

Is anyone in k12 algebra 1?

MsBrains (ms-brains):

Which grade are you in? I could help you if you have any questions.

OpenStudy (kimberlyevens):

I can message you because its block to people you didnt fan, 11th

OpenStudy (kimberlyevens):

@Ms-Brains

MsBrains (ms-brains):

Wow. Okay Let's start with question number 1. :)

MsBrains (ms-brains):

Any ideas?

MsBrains (ms-brains):

The prime factorization of 60 is: 60 = 2 * 2 * 3 * 5 This can also be written using exponents as: 60 = 2^2 * 3 * 5 Explanation 60 ÷ 2 = 30 30 ÷ 2 = 15 15 ÷ 3 = 5 5 ÷ 5 = 1